OpenStudy (anonymous):

If 75 equals a number plus four times the number, find the number A. -15 B. 30 C. 15 D. -30

OpenStudy (anonymous):

a + 4a = 75

OpenStudy (anonymous):

ah thanks. :)

OpenStudy (blurbendy):

5a = 75 a = 75 / 5 a = 15